What is the value today of $3,100 per year, at a discount rate of 9 percent, if the first payment is received 5 years from today and the last payment is received 15 years from today? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
1
Expert's answer
2020-04-30T09:54:26-0400

"PVA=C\times(\frac{1-\frac{1}{(1+r)^n}}{r})"

"PVA=3100\times(\frac{1-\frac{1}{(1+0.09)^{10}}}{0.09})"


PVA=19894.74


"PV=\frac{FV}{(1+r)^n}"


"PV=\frac{19894.74}{(1+0.09)^5}"


PV=12930.22